Use Cramer’s rule to solve
x+2y+z=5
2x+2y+z=6
x+2y+3z=9
1
Expert's answer
2020-03-05T16:55:36-0500

"D=\begin{vmatrix}
 1 & 2 & 1\\
 2 & 2 & 1\\
 1 & 2 &3
\end{vmatrix}=6+2+4-2-2-12=-4"

"D_x=\begin{vmatrix}
 5 & 2&1 \\
 6& 2&1\\
9&2&3
\end{vmatrix}=30+18+12-18-36-10=-4"

"D_y=\begin{vmatrix}
 1&5&1\\2&6&1\\1&9&3
\end{vmatrix}=18+18+5-6-30-9=-4"

"D_z=\begin{vmatrix}
 1&2&5\\2&2&6\\1&2&9
\end{vmatrix}=18+20+12-10-36-12=-8"

"x=D_x\/D=(-4)\/(-4)=1;"

"y=D_y\/D=(-4)\/(-4)=1;"

"z=D_z\/D=(-8)\/(-4)=2."

Answer:

"x=1; y=1; z=2."